OpenStudy (anonymous):

HELLLP!! Ill give medals..I'm working with probabilities!! Two number cubes, a red one and a blue one, are rolled. What is the probability that the red cube will be an odd number while the sum of the two dice is an even number?

OpenStudy (anonymous):

There are 4 cases on the sum of numbers 1.)odd+odd=even 2.)odd+even=odd 3.)even+odd=odd 4.)even+even=even now it is given that the sum is even thus we can remove case 2 and 3 the cases we have now are 1.)odd(red)+odd(blue)=even 2.)even(red)+even(blue)=even clearly probability that red will be odd is \[\frac{1}{2}\] Although I'm not sure myself

OpenStudy (anonymous):

I'll call someone to help you, @Kainui

OpenStudy (anonymous):

\[P(reddodd|sumeven)=\frac{P(redodd-sumeven )}{P(sumeven))}\] here by - I mean intersection, probability of red odd given that sum is even is given by probability of red odd and sum even divided by probabilityof sum even P(redodd-sumeven)=\[\frac{1}{4}\], which is probability of sum is even and red is odd \[P(sumeven)=\frac{2}{4}=\frac{1}{2}\] which is probability of sum is even \[P(redodd|sumeven)=\frac{\frac{1}{4}}{\frac{1}{2}}=\frac{1}{4} \times 2=\frac{1}{2}\]

OpenStudy (anonymous):

So I think it's \[\frac{1}{2}\] but you should confirm from somebody else

OpenStudy (anonymous):

@k_lynn @iGreen.

OpenStudy (mathmate):

"What is the probability that the red cube will be an odd number \(while\) the sum of the two dice \(is\) an even number?" I find the question not very explicit. The fact that the sum \(is\) an even number (and not "will be") seems to imply that it is a condition, in which case the conditional probability calculations done by @Nishant_Garg is good. There is a slight chance that the \(while\) is meant to be a joint condition, in which case we only have 1/4 of the possible cases. I.e. \(P(red~odd \cap sum~even)\)